# Minerva: A File-Based Ransomware Detector

### v2

@@ 6.3.1. Traditional Ransomware
Traditional ransomware comprises ransomware families that are used in real-world attacks and can be found in the wild. These families of ransomware typically exhibit low degrees of parallelism and spawn few processes to encrypt the target system and perform other malware-related activities, such as command-and-control communication (De Gaspari et al., 2020). Traditional ransomware sometimes employ multiple processes to speed up the encryption of the target system, but do not perform complex balancing of operations between the various processes to avoid detection (De Gaspari et al., 2022a).
Table 1 reports the performance of Minerva on the traditional ransomware test set. The true positive rate ($TPR_{T}$) value for all of the considered time windows is above $98\%$, reaching over $99\%$ on the $4sec$ time window. This means that all individual Minerva classifiers can accurately pinpoint traditional ransomware activity with less than 2% of false negatives. Similarly, the TNR ranges between ${\sim}98.5\%$ and ${\sim}99\%$ across all windows, highlighting the ability of Minerva to effectively detect ransomware without compromising user experience. Finally, the overall stacked performance of Minerva, which aggregates the predictions of each individual classifier to take a decision (see Section 4.1), reaches over $99\%$ TPR. The improved TPR of stacked Minerva compared to the $4sec$ window classifier derives from the independent nature of the errors made by each individual classifier, which results in higher overall detection rate.
The performance analysis in Table 1 emphasizes the ability of Minerva to generalize well across ransomware families, even when very short time windows are considered. This characteristic of Minerva is fundamental in ensuring rapid detection of ransomware activity ($0.52sec$ mean, see Section 6.2), which ensures low data loss rate and the compatibility of Minerva with near-zero overhead data protection techniques.
6.3.2. Evasive Multiprocess Ransomware
Table 1 reports the performance of Minerva on the evasive multiprocess ransomware test set. The test data is collected by running a variety of evasive techniques (De Gaspari et al., 2020), which are shown to completely evade behavioral-based ransomware detection (De Gaspari et al., 2022a). The true positive rate ($TPR_{M}$) is largely above $99\%$ for all of the time windows, which confirms our hypothesis that Minerva’s file-based behavioral profiling is resilient to this type of evasive multiprocess ransomware, contrary to prior works. We also note that there is essentially no performance variation between the different Minerva classifiers. While initially perplexing, this phenomenon can be explained upon closer analysis of the behavior of the multiprocess ransomware. By distributing the workload across multiple processes, the activity of the ransomware on each individual file typically lasts less than $0.5sec$ — the minimum time required for detection by Minerva with $K=2$. Consequently, all Minerva classifiers see the full ransomware activity on the file, which explains the negligible performance variations between different windows. Furthermore, as elaborated later in Section 6.6, the variability in behavior among the various processes of multiprocess ransomware is minimal compared to traditional ransomware, which aids in clarifying the near-perfect detection performance.
6.3.3. Comparison with SOTA
We further compare the performance of Minerva against well-known benchmark ransomware detection approaches (Continella et al., 2016; Mehnaz et al., 2018). Table 2 presents our comparison. As illustrated in the table, Minerva achieves performance comparable to state-of-the-art approaches against traditional ransomware. Furthermore, Minerva successfully detects evasive multiprocess ransomware, a feat that existing approaches fail to accomplish.
@@ 6.3.3. Comparison with SOTA
6.4. Unseen Detection Performance
Our prior results demonstrate that Minerva exhibits excellent performance in detecting known ransomware. Nevertheless, given the continual emergence of new ransomware variants and the ongoing evolution of existing ones, an arguably more critical evaluation criterion for ransomware detectors lies in their capability to generalize to unseen data.
In line with previous works (Continella et al., 2016; Mehnaz et al., 2018; Kharraz and Kirda, 2017), we assess Minerva’s capability to generalize to unseen ransomware by removing $10\%$ of the variants from our training data and evaluating detection performance on these withheld samples. This evaluation provides a measure of the expected predictive and generalization capabilities of Minerva on future, unknown ransomware variants. Table 3 reports the results of our test. As highlighted in the table, Minerva’s performance on unseen variants closely aligns with its performance on known variants, exhibiting only a minor performance degradation of ${\sim}0.2$ percentage points across all classifiers. These findings underscore the ability of Minerva to learn descriptive features that accurately capture the diverse range of behaviors displayed by ransomware.
@@ 6.5. Adaptive Detection Performance
As previously discussed, our primary objective with the design of Minerva is to develop a robust classifier, capable of withstanding adaptive attacks specifically engineered to evade its detection. As outlined in Section 4, we developed Minerva following a contrastive design, where altering one feature triggers changes to one or more other features. This characteristic renders evasion attempts by adaptive ransomware precarious, as any alteration intended to evade detection inadvertently induces changes in other features used by Minerva, ultimately leading to detection.
We assess the effectiveness of our contrastive design by evaluating Minerva against ransomware that has been specifically altered to evade Minerva detection mechanisms. This adaptive ransomware is an enhanced variant of the evasive multiprocess ransomware (De Gaspari et al., 2022a), improved to manipulate one or more of the features used by Minerva to circumvent detection, as elaborated in Section 5.2.2. We stress that all these adaptive families are not included in the training data, and are therefore unseen by the classifier. Table 4 illustrates the stacked detection rate of Minerva for different families of adaptive ransomware. Notably, the TPR remains largely unaffected by the behavioral modifications of the different adaptive families. Specifically, adaptive ransomware targeting the read-write entropy features are promptly detected by Minerva within $0.5sec$, which is the minimum detection time when $K=2$. Similarly, adaptive ransomware altering read- and write-ratio behaviors is also consistently detected. However, the mean detection time for this adaptive family is significantly extended (${\sim}3$x), with a $1sec$ standard deviation. These findings are expected, as this adaptive ransomware family reduces the R/W ratio feature by aggressively slowing down encryption activity (up to $80\%$ slowdown, as outlined in Section 5.2.2). Nevertheless, owing to the contrastive design of Minerva, detection performance remains excellent, as the classifiers leveraging long-term windows effectively identify the slow encryption activity. The sole adaptive family achieving limited success in evading Minerva are the adaptive operation number ransomware. These ransomware reduce the number of read and write operations by increasing the size of each individual operation. However, as highlighted in Table 4, Minerva continues to deliver excellent detection performance against this adaptive family. Furthermore, since this evasion strategy entails reducing the number of operations while increasing their size, attempting to evade Minerva using this method renders the attack easily detectable through process-based detection (De Gaspari et al., 2022a). Finally, we also evaluated Minerva against compatible combinations of evasive strategies (r/w entropy + r/w ratio, r/w ratio + op. num.), obtaining similar detection results.
@@
In Figure 5, we further study the detection rate distribution across Minerva’s classifiers for the different adaptive families. We exclude from this analysis r/w entropy adaptive family, as it is always detected by the short-term $0.25sec$ classifier. Figure 5(a) plots the detection rate against adaptive ransomware that manipulates the number of operations feature. From the figure, we notice that the vast majority of all the family’s variants ($>90\%$) are detected within $0.5sec$ by the short-term $0.25sec$ window when $K=2$. These results are expected, as this adaptive family does not attempt to slow down encryption activity, and therefore we expect a detection time in line with the overall mean detection time (see Figure 3). Figures 5(b) and 5(c) depict the detection rate distribution for the adaptive R/W ratio family of ransomware for $K=1$ and $K=2$ respectively. As expected, the detection distribution among the Minerva classifiers significantly differs compared to other ransomware families, with a tendency towards longer-term classifiers as the slowdown ratio increases. Furthermore, interesting insights can be derived by comparing the detection distribution for $K=1$ and $K=2$. Initially, one might expect the detection distribution for $K=2$ to closely resemble that of $K=1$, albeit with double the detection time. However, we observe that this expectation is incorrect. In reality, the minimum detection time for adaptive R/W ransomware increases fourfold between the two settings, while the maximum detection time increases twofold, as initially anticipated. This behavior is explained by the weaker detection performance of Minerva’s short-term classifiers on individual windows of ransomware activity for this adaptive family. While the short-term term classifiers effectively detect at least one window of malicious activity (sufficient for detection when $K=1$), they fail to detect two consecutive windows when the ransomware activity slows down. This ultimately leads to a bias in the detection distribution towards long-term classifiers, and, therefore, longer detection times.
@@ 6.6. Feature contribution to decision-making
Explainability is a complex yet crucial aspect in the deployment of DNNs, especially in critical domains such as cybersecurity. While previous sections have empirically showcased the effectiveness of Minerva, its ability to generalize to unseen data, and its resilience to adaptive evasion attacks, it is also important to provide insights into the reasons behind Minerva’s effectiveness. In this section, we aim to bridge this gap by examining the contributions of each of Minerva’s features to the final classification outcome. We leverage the SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) technique, a widely adopted method to interpret model predictions. SHAP values are based on a game theoretic approach to assign importance values to each feature in a model. Features with positive SHAP values influence the prediction towards the positive class, whereas those with negative SHAP values steer the prediction towards the negative class. The magnitude of each individual feature indicates the strength of its contribution to the final prediction. As a reminder, Minerva is a file-based behavioral classifier. Therefore, it is understood that the ensuing discussion on feature importance and behavioral profiles is conducted from the perspective of activity on individual files.
@@
Figure 6 presents the SHAP plots for the $1sec$ window Minerva classifier, highlighting the importance of each selected feature to the final classification and how feature importance changes for different types of ransomware. Each plot presents the SHAP values of Minerva’s features for 100 randomly sampled data points. For ransomware, we generate separate SHAP plots for each type, allowing us to highlight the distinctive contribution of features to classification, based on the diverse behavior exhibited by each ransomware type. Figures 6(a)6(b), and 6(c) display the individual contributions of the Minerva features to the prediction of benign, traditional ransomware and multiprocess ransomware classes for the one-second window model. An immediate observation is the significance of the write ratio feature as a distinguisher between benign and ransomware instances. This finding is expected, as benign applications typically perform partial writes to files, while ransomware aims to fully rewrite user files as part of its malicious activity. A similar pattern emerges with the read ratio feature, where benign applications often read only portions of files, while ransomware tend to read files fully to encrypt them. Figure 6(b), however, reveals exceptions to this trend, indicating that some traditional ransomware variants also exhibit partial file read and write behaviors. We hypothesize that these ransomware variants prioritize speed over completeness of encryption, sacrificing full file encryption to achieve higher overall encryption speed. Other important features are operation number and process number. Figure 6 underscores the descriptive nature of these features, as ransomware typically exhibits a notably higher number of operations and number of processes acting on files during the encryption activity, as previously highlighted in Figures 2(b) and 2(c). This characteristic is especially prominent for multiprocess ransomware, where multiple processes collaborate in the encryption process to mimic the per-process behavior of benign processes. Finally, the read and write entropy features play a crucial role in discriminating between benign and ransomware activity, with the write entropy feature having a particularly strong influence on the model’s classification for ransomware. Conversely, the impact of these features on benign activity is ambiguous. This outcome is expected, as certain benign applications display read/write entropy profiles akin to ransomware, as previously outlined in Section 4.2.
@@
Figures 6(e),and 6(f) illustrate the SHAP plots for adaptive R/W ratio and adaptive operation number, respectively. We observe that these plots exhibit minimal variation when compared to the non-adaptive multiprocess ransomware depicted in Figure 6(c). These findings are expected, as the behavioral changes introduced by these adaptive ransomware families have limited impact on individual features of Minerva. Specifically, for adaptive R/W ransomware, the behavioral changes typically affect only short-term classifiers, whose detection windows are too brief to capture a meaningful portion of the malicious behavior, resulting in classification failure. Conversely, for long-term classifiers, the entirety (or a significant portion) of the malicious activity is captured, leading to a behavioral profile — and hence a SHAP profile — closely resembling that of non-adaptive families. Likewise, adaptive operation number ransomware primarily alter the number of operations executed during encryption activities. This behavior is evident in Figure 6(f), where the contribution of the operation number feature on the final classification is negligible, whereas the contribution of the other features remains largely unchanged. Lastly, Figure 6(d) clearly illustrates the influence of contrastive features such as entropy and number of operations on classification outcome. We observe that while the entropy features exhibit a minimal to negative impact on ransomware classification, the contribution of the operation number features shows a considerable increase. This behavior is expected, as adaptive entropy ransomware relies on dummy, low entropy operations to decrease the average read/write entropy and evade detection.
At a higher level, Figures 6(a)6(b), and 6(c) provide insights on the detection performance of Minerva discussed in Section 6.3. Specifically, we observe that the SHAP plot for benign processes — and thus their behavioral profiles — display a greater degree of variability compared to ransomware, making it challenging for Minerva to learn a general representation for this class. Conversely, ransomware, particularly multiprocess ransomware, exhibit more uniform behavioral profiles, thereby simplifying the task of learning a suitable representation for classification.
6.7. Minerva’s detection module overhead
@@ 6.7.1. Disk activity monitor overhead
The Disk Activity Monitor module is a shim providing system-wide, real-time information on all file system operations performed on critical files (see Section 4.1.1). Minerva, similar to previous research in this domain (Continella et al., 2016; Mehnaz et al., 2018) relies on an I/O Request Packets Logger driver (IRPLogger) to capture file system activity of interest. This collected information is further processed to generate feature vectors utilized by Minerva for detection. Consequently, the overhead induced by the disk activity monitor primarily stems from the continuous operation of the IRPLogger. Previous in-depth analyses indicate that, on average, the IRPLogger incurs a $1.02\%$ increase in the system’s CPU usage (Mehnaz et al., 2018), with no notable performance degradation from the users’ standpoint (Continella et al., 2016). Additionally, it is important to highlight that Minerva maintains behavioral profiles exclusively for opened user files, further minimizing its impact on the system.

### v3

@@ 6.3.1. Traditional Ransomware
Traditional ransomware comprises ransomware families that are used in real-world attacks and can be found in the wild. These families of ransomware typically exhibit low degrees of parallelism and spawn few processes to encrypt the target system and perform other malware-related activities (De Gaspari et al., 2020). Traditional ransomware sometimes employ multiple processes to speed up the encryption of the target system, but do not perform complex balancing of operations between the various processes to avoid detection (De Gaspari et al., 2022a).
Table 1 reports the performance of Minerva on the traditional ransomware test set, which includes Traditional Classic ($TPR_{T}$) and Current ($TPR_{C}$). The true positive rates value for all the considered time windows is $\geq 98\%$, reaching over $99\%$ on the $4sec$ window. This means that all individual Minerva classifiers can accurately pinpoint traditional ransomware activity with less than 2% of false negatives. Similarly, the TNR ranges between $98\%$ and $99\%$ across all windows, highlighting the ability of Minerva to effectively detect ransomware without compromising user experience. Finally, the overall stacked performance of Minerva, which aggregates the predictions of each individual classifier to take a decision, reaches over $99\%$ TPR. The improved TPR of stacked Minerva compared to the $4sec$ window classifier derives from the independent nature of the errors made by each individual classifier, which results in higher overall detection rate. The performance analysis in Table 1 emphasizes the ability of Minerva to generalize well across ransomware families, including very recent (2022-25) samples. This characteristic of Minerva is fundamental in ensuring rapid detection of ransomware activity, discussed in Section 6.2, which ensures low data loss rate and the compatibility of Minerva with near-zero overhead data protection techniques.
@@
6.3.2. Evasive Multiprocess Ransomware
Table 1 reports the performance of Minerva on the evasive multiprocess ransomware test set. The true positive rate ($TPR_{M}$) is largely above $99\%$ for all of the time windows, which confirms our hypothesis that Minerva’s file-based behavioral profiling is resilient to this type of evasive multiprocess ransomware, contrary to prior works. We also note that there is essentially no performance variation between the different Minerva classifiers. While initially perplexing, this phenomenon can be explained upon closer analysis of the behavior of the multiprocess ransomware. By distributing the workload across multiple processes, the activity of the ransomware on each individual file typically lasts less than $0.5sec$ — the minimum time required for detection by Minerva with $K=2$. Consequently, all Minerva classifiers see the full ransomware activity on the file, which explains the negligible performance variations between different windows. Furthermore, as elaborated later in Section 6.5, the variability in behavior among the various processes of multiprocess ransomware is minimal compared to traditional ransomware, which aids in clarifying the near-perfect detection performance.
@@ 6.3.2. Evasive Multiprocess Ransomware
6.3.3. Unseen Detection Performance
Given the constant evolution of ransomware, an essential evaluation criterion for ransomware detectors is their ability to generalize to previously unseen data. We assess Minerva’s capability to generalize to unseen ransomware by randomly removing $10\%$ of the variants from our training data and evaluating detection performance on these withheld samples. This evaluation provides a measure of the expected predictive and generalization capabilities of Minerva on future, unknown ransomware variants. Table 2 reports the results of our test. As highlighted in the table, Minerva’s performance on unseen variants closely aligns with its performance on known variants, exhibiting only a minor performance degradation of ${\sim}0.2$ percentage points across all classifiers. These findings underscore the ability of Minerva to learn descriptive features that accurately capture the diverse range of behaviors displayed by ransomware.
@@ 6.4. Adaptive Detection Performance
We assess the effectiveness of Minerva’s contrastive design by evaluating its performance against three adaptive ransomware families specifically engineered to evade detection (see Section 5.2). Table 3 illustrates the stacked detection rate for these adaptive families. Notably, the TPR remains largely unaffected by the behavioral modifications of the different adaptive families. Specifically, adaptive ransomware targeting the read-write entropy features are promptly detected by Minerva within $0.5sec$, which is the minimum detection time when $K=2$. Similarly, adaptive ransomware altering read- and write-ratio behaviors are also consistently detected. However, the mean detection time for this adaptive family is significantly larger at $1.82\pm 1.07$sec. These findings are expected, as the reduction in the R/W ratio feature is achieved by aggressively slowing down encryption activity (up to $80\%$ slowdown, as outlined in Section 5.2). Nevertheless, detection performance remains excellent, as the classifiers leveraging long-term windows effectively identify the slow encryption activity. The sole adaptive family achieving limited success in evading Minerva are the adaptive operation number ransomware. These ransomware reduce the number of read and write operations by increasing the size of each individual operation. However, as highlighted in Table 3, Minerva continues to deliver excellent detection performance against this adaptive family. Furthermore, since this evasion strategy entails reducing the number of operations while increasing their size, attempting to evade Minerva using this method renders the attack easily detectable through process-based detection (De Gaspari et al., 2022a). Finally, we also evaluated Minerva against compatible combinations of evasive strategies (r/w entropy + r/w ratio, r/w ratio + op. num.), obtaining similar detection results.
@@
In Figure 5, we further study the detection rate distribution across Minerva’s classifiers for the different adaptive families. We exclude from this analysis r/w entropy adaptive family, as it is always detected by the short-term $0.25sec$ classifier. Figure 5(a) plots the detection rate against adaptive ransomware that manipulates the number of operations feature. From the figure, we notice that the vast majority of variants ($>90\%$) are detected within $0.5sec$ by the short-term $0.25sec$ window when $K=2$. These results are expected, as this adaptive family does not attempt to slow down encryption activity, and therefore we expect a detection time in line with the overall mean detection time. Figures 5(b) and 5(c) depict the detection rate distribution for the adaptive R/W ratio family of ransomware for $K=1$ and $K=2$ respectively. The detection distribution among the Minerva classifiers skews towards longer-term classifiers as the slowdown ratio increases, as expected. Furthermore, interesting insights can be derived by comparing the detection distribution for $K=1$ and $K=2$. Initially, one might expect the detection distribution for $K=2$ to closely resemble that of $K=1$, albeit with double the detection time. However, we observe that this expectation is incorrect. In reality, the minimum detection time for adaptive R/W ransomware increases fourfold between the two settings, while the maximum detection time increases twofold, as initially anticipated. This behavior is explained by the weaker detection performance of Minerva’s short-term classifiers on individual windows of ransomware activity for this adaptive family. While the short-term term classifiers effectively detect at least one window of malicious activity (sufficient for detection when $K=1$), they fail to detect two consecutive windows when the ransomware activity slows down. This ultimately leads to a bias in the detection distribution towards long-term classifiers, and, therefore, longer detection times.
@@ 6.5. Feature contribution to decision-making
This section assesses the explainability of Minerva and the contributions of each of feature to the final classification outcome. We leverage the SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) technique, a widely adopted method to interpret model predictions. Features with positive SHAP values influence the prediction towards the positive class (and vice-versa), while the magnitude of each individual feature indicates the strength of its contribution to the final prediction. As a reminder, Minerva is a file-based behavioral classifier. Therefore, it is understood that the ensuing discussion on feature importance and behavioral profiles is conducted from the perspective of activity on individual files.
@@
Figure 6 presents the SHAP plots for the $1sec$ window Minerva classifier, highlighting the importance of each selected feature to the final classification and how feature importance changes for different types of ransomware. Each plot presents the SHAP values of Minerva’s features for 100 randomly sampled data points. Figures 6(a)6(b), and 6(c) display the individual contributions of the Minerva features to the prediction of benign, traditional ransomware and multiprocess ransomware classes for the one-second window model. An immediate observation is the significance of the write ratio feature as a distinguisher between benign and ransomware. This finding is expected, as benign applications typically perform partial writes to files, while ransomware aims to fully rewrite user files as part of its malicious activity. A similar pattern emerges with the read ratio feature, where benign applications often read only portions of files, while ransomware tend to read files fully to encrypt them. Figure 6(b), however, reveals exceptions to this trend, indicating that some traditional ransomware variants also exhibit partial file read and write behaviors. We hypothesize that these ransomware variants prioritize speed over completeness of encryption, sacrificing full file encryption to achieve higher overall encryption speed. Other important features are operation number and process number, with ransomware typically exhibiting a notably higher number of operations and processes acting on files during the encryption activity. This characteristic is especially prominent for multiprocess ransomware. Finally, the read and write entropy features play a crucial role in detection, with the write entropy feature having a particularly strong influence on the model’s classification for ransomware. Conversely, the impact of these features on benign activity is ambiguous. This outcome is expected, as certain benign applications display read/write entropy profiles akin to ransomware, as previously outlined in Section 4.2.
@@
Figures 6(e) and 6(f) illustrate the SHAP plots for adaptive R/W ratio and adaptive operation number, respectively. We observe that these plots exhibit minimal variation when compared to the non-adaptive multiprocess ransomware depicted in Figure 6(c). These findings are expected, as the behavioral changes introduced by these adaptive ransomware families have limited impact on individual features of Minerva. For adaptive R/W ransomware, the behavioral changes typically affect only short-term classifiers, while long-term classifiers still capture the complete malicious activity. Adaptive operation number ransomware primarily alter the number of operations executed during encryption activities, which is evident in Figure 6(f), where the contribution of the operation number feature on the final classification is negligible. However, the remaining features still lead to detection. Lastly, Figure 6(d) clearly illustrates the influence of contrastive features such as entropy and number of operations on classification outcome. We observe that while the entropy features exhibit a minimal to negative impact on ransomware classification, the contribution of the operation number features shows a considerable increase. This behavior is expected, as adaptive entropy ransomware relies on dummy, low entropy operations to decrease the average read/write entropy and evade detection.
(a) Overhead on file read operations under continuous I/O load.
(b) Overhead on file write operations under continuous I/O load.
(c) Memory overhead under different I/O conditions. Rolling averages as dashed lines.
Figure 7. Overall system overhead introduced by Minerva.
6.6. Overhead Analysis
@@ 6.6.1. Disk Activity Monitor overhead
Minerva’s Disk Activity Monitor module leverages an I/O Request Packets Logger driver (IRPLogger) to capture relevant file system activity (Continella et al., 2016). To evaluate the overhead introduced by the Disk Activity Monitor, we perform continuous read and write operations of different sizes ranging from 1KB to 100MB, simulating workloads on both small and large files, and measure the corresponding completion times. Each operation was repeated $1,000$ times for reading and $1,000$ times for writing for each considered operation size, resulting in a total of $28,000$ file system operations. To avoid caching-related side effects, all operations were conducted on separate individual files. The tests were performed on a Windows 11 virtual machine equipped with an NVMe solid-state drive, 16 CPU cores, and 16GB of RAM. As shown in Figures 7(a) and 7(b), Minerva’s overhead during continuous I/O operations remains low, ranging from $0-20\%$ for all read and write operations, with an outlier reaching $\sim 35\%$. On average, the overhead recorded across all operations was $6.76\%(\pm 11.49)$.
To further characterize the resource requirements of Minerva, we analyze its main memory and CPU usage under varying I/O loads. Figure 7(c) presents Minerva’s main memory usage in KB as processes perform operations at random intervals, ranging from a few seconds for low I/O to continuous operations at full I/O. Memory usage exhibits a consistent pattern across all I/O loads, with rapid increases in the first few seconds of activity followed by fluctuations based on the number of requests and the size of the operations. The running average, represented by the dashed lines in the figure, stabilizes early in the execution and remains relatively constant throughout. This behavior aligns with Minerva’s design, as it only stores file system operations within a maximum time window of 4 seconds, ensuring efficient memory usage and scalability even under high-intensity I/O workloads. Across our tests, we observe a negligible increase of $2.31(\pm 1.09)$ percentage points in CPU usage when Minerva is active in the system.
Finally, we ran a small scale user study to assess whether there is any perceived usability degradation when using Minerva in day-to-day tasks. We asked 5 users to interact with two systems VMs: one running Minerva and one without. The users were allowed to freely use the system, but were asked to include browsing, document editing, software installation, and file transfers operations. After they completed their activity, we asked them to evaluate their experience on a scale from 1 to 5 with each system, where 1 corresponded to “unusable system” and 5 corresponded to “perfectly usable”. Users reported an average of $4.8$ and $4.6$ for the system with and without Minerva, respectively. We attribute the perceived usability degradation on the system without Minerva to the use of virtualization.